# Online Cascade Learning for Efficient Inference over Streams

### v1

@@ Abstract
Large Language Models (LLMs) have a natural role in answering complex queries about data streams, but the high computational cost of LLM inference makes them infeasible in many such tasks. We propose online cascade learning, the first approach to addressing this challenge. The objective here is to learn a “cascade” of models, starting with lower-capacity models (such as logistic regressors) and ending with a powerful LLM, along with a deferral policy that determines the model that is used on a given input. We formulate the task of learning cascades online as an imitation-learning problem and give a no-regret algorithm for the problem. Experimental results across four benchmarks show that our method parallels LLMs in accuracy while cutting down inference costs by as much as 90%, underscoring its efficacy and adaptability in stream processing.
@@ 1 Introduction
Large language models (LLMs) (Bommasani et al., 2021; Touvron et al., 2023b; Brown et al., 2020) hold great promise as a means of one-pass query answering over streams of text documents. For example, suppose we have a stream of movie reviews posted on the internet and would like to retrieve the ones that are positive (Figure 1). To solve such a problem, we could create a prompt that asks the LLM to determine the sentiment behind a review. The LLM could then use this prompt to process each statement in sequence.
However, a fundamental difficulty here is that inference using LLMs is expensive. Based on our benchmarking (Appendix B.1), it takes 3.6 seconds for a GPU server with eight A100 GPUs to process a document having 8,192 tokens using the largest Llama (Touvron et al., 2023a). To process one million such documents per hour would thus require 1000 A100 servers. Using Amazon Web Services, such a computation would cost more than 30,000 USD per hour, if it were even possible to procure the machines required.
Figure 1: A sentiment analysis task over a stream of movie reviews from IMDB (Maas et al., 2011). Our method uses the cheapest logistic regression model (green lines) to process the simpler inputs (reviews) and defers more complex inputs to the larger models (orange and red lines). When the cascade proceeds to the LLM, the annotations are collected to update the smaller models online.
There are two popular ways to reduce the cost of LLM inference. The first is to distill a large language model into a smaller model that can process a document using less computation (Hinton et al., 2015; Gu et al., 2023; Hsieh et al., 2023). The second is to use a cascade of models that use smaller models to process “easier” inputs and reserve the largest models for the most difficult inputs (Varshney & Baral, 2022; Chen et al., 2023). However, these existing proposals assume a model of learning in which a labeled training set is available beforehand, making them unsuitable for a streaming setting. By contrast, the streaming setting requires online learning (Hoi et al., 2021).
@@
In this paper, we propose online cascade learning as a way to bridge this gap in the literature. Our goal here is to learn a cascade of models, arrayed from least expensive to most, that is used to process the documents in a data stream in a single pass given a natural language prompt. The critical difference from prior work is that our cascades are trained in a fully online fashion, and require no pre-training, and no human-labeled training data.
A key component of our cascades is a deferral policy that decides, given an input, the “level” of the cascade that should handle the input (Figure 2). At startup, the policy keeps its “gates” open, so that the initial inputs flow through the cascade and are ultimately processed using the most expensive model (an LLM). These, in turn, become labeled training examples that are fed back into the cascade to train the less expensive models, as well as the deferral policy. Over time, as the model sees more data, the system stabilizes at a fixed point where the smaller, less expensive models process most of the data. The framework includes a hyperparameter that guides the trade-off between accuracy and cost.
We formalize the problem of online cascade learning in terms of an episodic Markov decision process (MDP) that takes both prediction loss and computational costs into consideration. We assume an expert policy — a high-capacity LLM — for this MDP. We learn the various components of the cascade through imitation learning (Ross et al., 2011) based on this expert. We show that our algorithm comes with a theoretical no-regret guarantee. Our experimental results, on four tasks of various complexity, show that our cascades can achieve accuracy comparable to LLM at a vastly reduced inference cost.
@@
  • We introduce online cascade learning, a new framework for learning model cascades for resource-intensive streaming analytics tasks. The framework allows for systematic tradeoffs between prediction accuracy and resource usage while allowing learning without any human-provided training data.
@@
  • We present rigorous experiments that show that our learning algorithm can achieve comparable accuracy as LLMs while saving up to $90\%$ of the inference costs. The source code is available at https://anonymous.4open.science/r/online_cascade_learning.
@@ 2.1 Preliminaries
Consider a set of $N$ models, $(m_{1},...,m_{N})$, arranged in ascending order of computational cost, and an evaluation dataset $X$ consisting of instances $x_{t}$. Each instance $x_{t}$ represents a query or input instance. For every query $x_{t}$, the cascade system aims to output a prediction $y_{t}$ from a label set $Y$. The computational cost of using a model $m_{i}$, which is assumed to be independent of $x_{t}$, is denoted as $c_{i}$. Thus, in a cascade setting, the model costs monotonically increase as $c_{1}<c_{2}<...<c_{N}$.
@@
    • exit resets the MDP with a new $x_{t+1}\in X$.
@@ 2.2 Online Cascade Learning
Therefore, the online learning of a cascade is equivalent to solving the above episodic MDP where the maximum length of each episode is $N$ – the number of models in our cascade. Within this MDP, both states and actions are delineated to be independent across different episodes.
The primary objective of online cascade learning is to minimize the cumulative cost across all conceivable states over a time horizon of $T$ episodes. This cost function should encompass two critical aspects: model prediction loss and computation costs. Thus, minimizing the cumulative cost is inherently a task of achieving a delicate balance between enhancing predictive accuracy and ensuring cost efficiency.
Formally, we use $\Pi$ to denote the complete policy space under consideration for the entire cascade. Within this space, for a given policy $\pi$, the notation $d_{\pi}$ is employed to signify the expected state distribution when the policy $\pi$ is applied consistently throughout an entire episode. The total cost of executing policy $\pi$ for $T$ queries can be represented as
$J(\pi)$ $=\sum^{T}_{t=1}\mathbb{E}_{s\sim d_{\pi}}[C_{\pi}(s)]$ (1)
@@
Since the query set $X$ is assumed to be IID and the states among episodes are independent, we can simplify the notation of state $\langle x_{t},i\rangle=s_{i}$. Let $p(\pi,s_{i})^{\prime}$ represents the probability of choosing an action $a_{i}=\texttt{defer}$ at $s_{i}$, and $1-p(\pi,s_{i})^{\prime}$ represents the probability of choosing an action $a_{i}\in Y$. The expected immediate cost at state $s_{i}$ can be written as:
@@
where $\mathcal{L}(a_{i}|x_{t})$ represents the prediction loss of the underlying model $m_{i}$ for query $x_{t}$ if the policy chooses to output a prediction at the current state, and $\mu c_{i+1}$ indicates the weighted computational costs to be incurred if the policy defers to the next model in the cascade.
Moreover, within a single episode, it is essential to accumulate the expected costs across all potential states. This necessitates accounting for the probability of a policy $\pi$ reaching a specific state $s_{i}$, denoted as $p^{s_{i}}_{\pi}$. Since a state $s_{i}$ can be reached if and only if the policy chooses to defer in all preceding states $s_{1},...,s_{i-1}$ within the same episode, therefore for a policy $\pi$, $p^{s_{i}}_{\pi}=\prod^{i-1}_{j=1}p(\pi,s_{j})^{\prime}$.
@@
  Initialize $\mathcal{D}\leftarrow\emptyset$
@@
        Let $\pi_{t}=\beta_{t}\pi^{*}+(1-\beta_{t})\hat{\pi}_{t}$ by setting   $\pi_{t}\leftarrow\pi^{*}$ at a probability $\beta_{t}$
@@ 3 Methodology
The overall process of the algorithm can be summarized as follows. At the initial iterations (i.e., the warm-up stage), it always defers to use the expert’s policy (i.e., the LLM) to gather an annotated dataset, denoted as $\mathcal{D}$. Then, it trains a new policy $\hat{\pi}$ to mimic the expert on the trajectories in $\mathcal{D}$. At iteration $t$, the algorithm collects more trajectories using $\hat{\pi}_{t}$ and obtains expert policy annotation at a decaying probability $\beta_{t}\neq 0$, in order to better leverage the presence of the expert policy and improve robustness.
At a high level, in the streaming setup, the online cascade learning algorithm continuously collects annotations from the expert policy (e.g., at a decaying probability $\beta_{t}$ or when the state reaches $s_{N}$ in the episode) and updates the policy through online gradient descent based on previous annotations. The general algorithm is detailed in Algorithm 1.
@@ 3.1 Theoretical Analysis
In online learning, a policy’s regret over time $T$ is its total cost minus the cost of the best fixed expert policy in hindsight. For online cascade learning, the regret of a learned policy $\pi$ can be defined as:
@@ 3.2 Confidence Calibration
In our work, since inference is performed in a zero-shot setting, we use the predictions of the LLM (i.e., $\operatorname*{arg\,max}m_{N}(x_{t})$) as the approximation of the ground truth labels $y^{*}_{t}$. During the calibration, the learning of $\pi_{i}^{\prime}$ would only optimize the parameters of the MLP, not the models $m_{i}$ in the cascade. During inference time, the post-hoc deferral policy would choose to defer if $\pi^{\prime}_{i}(m_{i}(x_{t}))>0.5$. Otherwise, it would output the prediction of the current model.
@@
Online Ensemble Learning 86.73 88.80 89.95 82.61 |||| 76.75 76.43 |||| 85.23 81.55 |||| 80.30 56.56 60.42 61.78 61.69 69.78 76.67
Online Cascade Learning 87.95 92.48 93.01 82.66 |normal-||| 82.36 82.75 |||| 82.36 83.26 |||| 81.03 60.78 65.34 69.75 61.95 71.86 78.49
@@
Table 1: Comparison of accuracy (and recall for HateSpeech dataset) among different methods under various cost budgets. The upper part of the table uses GPT-3.5 Turbo as the LLM in the cascade, while the lower part employs Llama 2 70B Chat. To ensure fairness, the same annotation cost budgets (i.e., the maximum allowable LLM calls, denoted as $\mathcal{N}$, controlled via adjusting the cost weighting factor $\mu$ for online cascade learning) are applied across all methods.
@@ 4 Experimental Setup
In our experiments, we construct a cascade system using three models: (i) a logistic regression model, (ii) a pretrained BERT-base model (Kenton & Toutanova, 2019), and (iii) GPT-3.5 Turbo [^1] https://platform.openai.com/docs/models/gpt-3-5. This diverse set of models allows us to evaluate the effectiveness and robustness of our online cascade learning approach across various complexities and types of queries.
To further test the adaptability and scalability of our system, we also conduct supplementary experiments where (a) Llama 2 70B Chat (Touvron et al., 2023b) is used as the alternative LLM, and (b) BERT-large is incorporated to create a larger cascade. These variations aim to demonstrate the flexibility of our framework in accommodating different cascade sizes and structures.
@@ 4.2 Baselines
  • LLMs in the Cascade: This includes GPT-3.5 Turbo and Llama 2 70B Chat. The prompts used for zero-shot task instructions are outlined in Appendix B.2.
  • Knowledge Distillation: We fine-tune smaller models using a portion of LLM annotations. To ensure fairness, datasets are split equally, with 50% allocated for training (as distillation labels) and the remaining 50% for testing. All methods are evaluated using these identical test sets.
@@ 5.1 Overall Performance and Cost Trade-offs
The results on the HateSpeech dataset further reveal the strengths of our online cascade learning system in handling datasets with significant class imbalance. Most models may face a trade-off between accuracy and recall due to the imbalanced nature of HateSpeech. However, the accuracy-cost and recall-cost trade-offs, respectively depicted in the upper and lower subplots of Figure 3 and 4, demonstrate that our system effectively improves recall with minimal impact on accuracy as the cost budget increases. Although the accuracy or recall of the online cascade learning might be marginally lower than the baselines in some cases, it consistently achieves a better balance between these two metrics. In particular, as demonstrated in Table 1 where Llama 2 70B Chat is the LLM and the cost budget $\mathcal{N}=4900$, our system even outperforms the LLM with a similar recall (ours: 82.03% vs. LLM: 82.19%) and a better accuracy (78.32% vs. 77.81%), underlining its effectiveness in handling imbalanced data streams.
@@
FEVER is a significantly more complex dataset compared to the previous benchmarks. It demands models to reason over the statements and validate their factuality based on parametric knowledge. Therefore, small models of limited capacities, such as logistic regression, struggle to perform effectively on FEVER even after several iterations of update, as evident in Table 1 where distilled LR can perform only slightly better than random guess (i.e., 50%). Recognizing these limitations, our online cascade learning system smartly adapts by prioritizing more capable models, such as BERT-base and LLM, for processing most of the queries, leading to a favorable accuracy-cost trade-off. Remarkably, when using Llama 2 70B Chat as the LLM with a cost budget of $\mathcal{N}=2800$, our system slightly outperforms the LLM in accuracy (77.73% vs. 77.15%), showcasing the system’s proficiency at navigating intricate reasoning tasks with enhanced cost-efficiency.

### v2

@@ Abstract
Large Language Models (LLMs) have a natural role in answering complex queries about data streams, but the high computational cost of LLM inference makes them infeasible in many such tasks. We propose online cascade learning, the first approach to address this challenge. The objective here is to learn a “cascade” of models, starting with lower-capacity models (such as logistic regression) and ending with a powerful LLM, along with a deferral policy that determines the model to be used on a given input. We formulate the task of learning cascades online as an imitation-learning problem, where smaller models are updated over time imitating the LLM expert demonstrations, and give a no-regret algorithm for the problem. Experimental results across four benchmarks show that our method parallels LLMs in accuracy while cutting down inference costs by as much as 90% with strong robustness against input distribution shifts, underscoring its efficacy and adaptability in stream processing.
@@ 1 Introduction
Large language models (LLMs) (Bommasani et al., 2021; Touvron et al., 2023b; Brown et al., 2020) hold great promise as a means of one-pass query answering over text streams. For example, suppose we have a stream of movie reviews posted on the internet and would like to retrieve the ones that are positive (Figure 1). To address this without human annotations, we can create a prompt for the LLM to identify the sentiment in a review. The LLM can then use this prompt to process each statement in sequence.
However, LLM inference can be extremely expensive, especially in a streaming setup where queries arrive continuously. Based on our benchmarking (Appendix B.1), a GPU server with eight A100 GPUs takes 3.6 seconds to process a document containing 8,192 tokens using the largest Llama (Touvron et al., 2023a). To process one million such documents per hour would thus require 1,000 A100 servers. Using Amazon Web Services, this computation would cost more than 30,000 USD per hour, if it were even possible to procure the machines required.
Figure 1: A sentiment analysis task over a stream of IMDB movie reviews (Maas et al., 2011). We use the cheapest logistic regression model (green lines) to process simpler queries and defer more complex queries to the larger models (orange & red lines). When the cascade proceeds to the LLM, the annotations are collected to update the smaller models online (blue lines).
There are two popular ways to reduce the cost of LLM inference. The first is to distill a large language model into a smaller model that can process a document using less computation (Hinton et al., 2015; Gu et al., 2023; Hsieh et al., 2023). The second is to use a cascade of models that use smaller models to process “easier” inputs and reserve the largest models for the most difficult inputs (Varshney & Baral, 2022; Chen et al., 2023). However, these existing proposals assume a model of learning in which a labeled training set is available beforehand, making them unsuitable for a streaming setting. In contrast, the streaming setting requires online learning (Hoi et al., 2021).
@@
In this paper, we propose online cascade learning to bridge this gap in the literature. Our goal here is to develop a cascade of models, arrayed from the least to the most complex, that can process queries in a data stream with optimal cost-performance trade-offs. The critical difference from prior work is that our cascades are trained in a fully online manner and do not require any human-labeled training data. The smaller models in the cascade would continuously evolve and improve over time by imitating the LLM’s demonstrations on the harder queries, enabling them to handle an expanding range of queries with increasing proficiency.
A key component of our cascades is the deferral policy that decides, given an input, the best “level” of the cascade that should handle the input (Figure 2). At startup, the policy keeps its “gates” open, allowing all initial inputs to flow through the cascade and be processed by the most expensive model (an LLM). These processed inputs then become training labels for updating the smaller models and the deferral policy within the cascade. Over time, as the model sees more data, the system stabilizes at a state where the smaller, less expensive models can handle the majority of the new inputs. The framework also incorporates a set of learning hyperparameters that adjust the trade-off between accuracy and cost based on user needs.
We formalize the problem of online cascade learning in terms of an episodic Markov decision process (MDP) that considers both prediction loss and computational costs for co-optimization. We assume an expert policy — a high-capacity LLM — for this MDP. We learn the various components of the cascade through imitation learning (Ross et al., 2011) based on the LLM demonstrations. We show that our algorithm comes with a theoretical no-regret guarantee. Our experimental results, on four tasks of various complexity, show that our proposed method can achieve accuracy comparable to LLM at a vastly reduced inference cost.
@@
  • We introduce online cascade learning, a new framework for learning model cascades in resource-intensive streaming analytics settings. The framework enables systematic trade-offs between prediction accuracy and resource usage, and allows learning without any human annotations.
@@
  • We present rigorous experiments showing that our proposed algorithm can achieve comparable accuracy as LLMs while saving up to $90\%$ of the inference costs. Our source code is available at https://github.com/flitternie/online_cascade_learning.
@@ 2.1 Preliminaries
Consider a set of $N$ models, $(m_{1},...,m_{N})$, arranged in ascending order of computational cost, and an evaluation data stream $X$ consisting of user queries $x_{t}$ that arrives at time $t$. For every query $x_{t}$, the cascade system aims to output a prediction $y_{t}$ from a label set $Y$. The inference cost of a model $m_{i}$, which is assumed to be independent of $x_{t}$, is denoted as $c_{i}$. In a cascade setting, the model costs monotonically increase as $c_{1}<c_{2}<...<c_{N}$.
@@
    • exit ends the current episode and resets the MDP with $\langle x_{t+1},1\rangle$ as the initial state of a new episode.
@@ 2.2 Online Cascade Learning
Therefore, the online learning of a cascade is equivalent to solving the above episodic MDP, where the maximum length of each episode is $N$, the number of models in the cascade. Within this MDP, both states and actions are delineated to be independent across different episodes and input queries.
The primary objective of online cascade learning is to minimize the cumulative cost across all conceivable states over a time horizon of $T$ episodes. This cost function should encompass two critical aspects: model prediction loss and computational costs. Thus, minimizing the cumulative cost is inherently a task of achieving a delicate balance between enhancing predictive accuracy and ensuring cost efficiency.
Formally, we use $\Pi$ to denote the complete policy space under consideration, which is the theoretical aggregation of all trainable model weights in the cascade. Within this space, for a given policy $\pi$, i.e., a cascade of models of certain weights, the notation $d_{\pi}$ is employed to signify the expected state distribution when the policy $\pi$ is applied consistently throughout an entire episode. The total cost of executing policy $\pi$ for $T$ episodes can be represented as:
$J(\pi)$ $=\sum^{T}_{t=1}\mathbb{E}_{s\sim d_{\pi}}[C_{\pi}(s)],$ (1)
@@
Since the query stream $X$ is assumed to be IID and the states among episodes are independent, we can simplify the notation of state $\langle x_{t},i\rangle$ as $s_{i}$. Let $p(\pi,s_{i})^{\prime}$ represents the probability of choosing an action $a_{i}=\texttt{defer}$ at $s_{i}$, and $1-p(\pi,s_{i})^{\prime}$ represents the probability of choosing an action $a_{i}\in Y$. The expected immediate cost at state $s_{i}$ can be written as:
@@
where $\mathcal{L}(a_{i}|x_{t})$ represents the prediction loss for query $x_{t}$ if the policy chooses to output a prediction at the current state using model $m_{i}$, and $\mu c_{i+1}$ indicates the weighted computational costs to be incurred if the policy defers to the next model $m_{i+1}$ in the cascade.
Moreover, within a single episode, it is essential to accumulate the expected costs across all potential states. This necessitates accounting for the probability of a policy $\pi$ reaching a specific state $s_{i}$, denoted as $p^{s_{i}}_{\pi}$. Since a state $s_{i}$ can be reached if and only if the policy chooses to defer in all preceding states $s_{1},...,s_{i-1}$ within the current episode, therefore for a policy $\pi$, $p^{s_{i}}_{\pi}=\prod^{i-1}_{j=1}p(\pi,s_{j})^{\prime}$. Please note that, the probability distribution over these states represents potential transitions, not an aggregation of model weights.
@@
  Initialize $\mathcal{D}\leftarrow\emptyset$, policy weights $\hat{\pi}_{1}$, and $\beta_{1}$
@@
        Let $\pi_{t}=\beta_{t}\pi^{*}+(1-\beta_{t})\hat{\pi}_{t}$ by setting   $\pi_{t}\leftarrow\pi^{*}$ at a decaying probability $\beta_{t}$
@@ 3 Methodology
The overall process of the algorithm is detailed in Algorithm 1, where $\pi^{*}$ refers to the LLM expert for imitation, and $\hat{\pi}_{1}$ indicates the initial states of the small models. In practice, the models are initialized either randomly or with their respective pretrained weights. For every incoming query $x_{t}$ at episode $t$ (outer loop), we sequentially utilize the $i$-th model in the current cascade to make a prediction (inner loop). Specifically, the model outputs a class probability distribution, which is converted by the policy into a deferral probability score to determine the action $a^{i}_{t}=\pi_{t}(s_{i})$
If the action is defer, the query would be navigated to the cascade’s next $(i+1)$-th model. Otherwise, if the action is to output a prediction at the current cascade level, it will make a prediction $y_{t}\leftarrow a^{i}_{t}$, and break the inner loop (succeeding models will not be activated for current query $x_{t}$). If the $N$-th model in the cascade is reached, that means the query has been deferred to the LLM expert $\pi^{*}$, thereby we collect its annotation $\{x_{t},y_{t}\}$ and aggregate it to the dataset $\mathcal{D}$.
By default, $\pi_{t}(s_{i})$ invokes the $i$-th model $\hat{\pi}_{t}(s_{i})=m_{i}^{t}(x_{t})$ for making an action at episode $t$. However, throughout the decision-making process, the cascade may optionally skip the rest layers and directly jump to the LLM expert $\pi^{*}$ at a non-zero decaying probability $\beta_{t}$ to obtain its demonstrations $\{x_{t},y_{t}\}$, similar to DAgger (Ross et al., 2011), for faster convergence. Therefore, in Algorithm 1, $\pi_{t}=\beta_{t}\pi^{*}+(1-\beta_{t})\hat{\pi}_{t}$ represents an expected value of the acting policy, not the combination of model parameters.
After processing each query, it outputs the prediction $y_{t}$, then update the policy $\hat{\pi}_{t+1}$, including the small models and deferral policy, to mimic the LLM expert on the collected trajectories $\mathcal{D}$. At a high level, the algorithm continuously collects annotations from the expert (e.g., at a decaying probability $\beta_{t}$ or when the state reaches $s_{N}$ in the episode) and updates the policy through online gradient descent based on expert demonstrations. Practically, the user can change the cost weighting factor $\mu$ in the objective function $J(\pi)$ and the initial decaying factor $\beta_{1}$ for adjusting cost budgets.
@@ 3.1 Theoretical Analysis
In online learning, a policy’s regret over time $T$ is its total cost minus the cost of the best fixed policy in hindsight. For online cascade learning, the regret of a learned policy $\pi$ can be defined as:
@@ 3.2 Confidence Calibration
In our work, since inference is performed in a zero-shot setting without any human annotations, we use the predictions of the LLM (i.e., $\operatorname*{arg\,max}m_{N}(x_{t})$) as the approximation of the ground truth labels $y^{*}_{t}$. Thus, calibration is only performed on those input queries where the LLM is invoked. During the calibration, the learning of $\pi_{i}^{\prime}$ would only optimize the parameters of the MLP, not the models $m_{i}$ in the cascade. During inference time, the post-hoc deferral policy would choose to defer if $\pi^{\prime}_{i}(m_{i}(x_{t}))>0.5$. Otherwise, it would output the prediction of the current model.
@@
Online Ensemble Learning 86.73 88.80 89.95 82.61 |||| 76.75 77.48 |||| 76.89 81.55 |||| 80.30 56.56 60.42 61.78 61.69 69.78 76.67
Online Cascade Learning 87.95 92.48 93.01 82.66 |||| 82.36 85.35 |||| 77.20 83.26 |||| 81.03 60.78 65.34 69.75 61.95 71.86 78.49
@@
Table 1: Comparison of accuracy (and recall for HateSpeech dataset) among different methods under various cost budgets. The upper part of the table uses GPT-3.5 Turbo as the LLM in the cascade, while the lower part employs Llama 2 70B Chat. To ensure fairness, the same annotation cost budgets (i.e., the maximum allowable LLM calls, denoted as $\mathcal{N}$, controlled via adjusting the cost weighting factor $\mu$ and decaying factor $\beta$ for online cascade learning) are applied across all methods.
@@ 4 Experimental Setup
In our experiments, we construct a cascade system using three models: (i) a logistic regression model, (ii) a pretrained BERT-base model with 110M parameters (Kenton & Toutanova, 2019), and (iii) GPT-3.5 Turbo [^1] https://platform.openai.com/docs/models/gpt-3-5. This diverse set of models allows us to evaluate the effectiveness and robustness of our online cascade learning approach across various complexities and types of queries.
To further test the adaptability and scalability of our system, we also conduct supplementary experiments where (a) a Llama 2 70B Chat (Touvron et al., 2023b) is used as the alternative LLM, and (b) a BERT-large with 340M parameters is incorporated to create a larger cascade. These variations aim to demonstrate the flexibility of our framework in accommodating different cascade sizes and structures.
@@ 4.2 Baselines
  • LLMs in the Cascade: This includes GPT-3.5 Turbo and Llama 2 70B Chat with zero-shot task prompting (details in Appendix B.2). The LLM outputs are also used as the annotations for online cascade learning and distillation.
  • Knowledge Distillation: We fine-tune smaller models using different portions of LLM annotations. To ensure fairness, datasets are split equally, with 50% prepared for training (as distillation labels) and the remaining 50% for testing. All methods are evaluated on the identical test sets. In our experiments, the distilled smaller models are used in isolation without any ensemble or cascade.
@@ 5.1 Overall Performance and Cost Trade-offs
The results on the HateSpeech dataset further reveal the strengths of our online cascade learning system in handling datasets with significant class imbalance. Most models may face a trade-off between accuracy and recall due to the imbalanced nature of HateSpeech. However, the accuracy-cost and recall-cost trade-offs, respectively depicted in the upper and lower subplots of Figure 3 and 4, demonstrate that our system effectively improves recall with minimal impact on accuracy as the cost budget increases. Although the recall rate of online cascade learning is marginally lower than the baselines under certain budgets, it can achieve a better balance between recall and precision, as evidenced by its consistently higher F1 scores in Appendix Figure 10 and 11. In particular, as demonstrated in Table 1 where Llama 2 70B Chat is the LLM and the cost budget $\mathcal{N}=4900$, our system even outperforms the LLM with a similar recall (ours: 82.03% vs. LLM: 82.19%) and a better accuracy (78.32% vs. 77.81%), underlining its effectiveness in handling imbalanced data streams.
@@
FEVER is a significantly more complex dataset compared to the previous benchmarks. It demands models to reason over the statements and validate their factuality based on parametric knowledge. Therefore, small models of limited capacities, such as logistic regression, struggle to perform effectively on FEVER even after several iterations of update, as evident in Table 1 where distilled LR can perform only slightly better than random guess (i.e., 50%). Recognizing these limitations, our online cascade learning system smartly adapts by prioritizing more capable models, such as BERT-base and the LLM, for processing most of the queries, leading to a favorable accuracy-cost trade-off. Remarkably, when using Llama 2 70B Chat as the LLM with a cost budget of $\mathcal{N}=2800$, our system slightly outperforms the LLM in accuracy (77.73% vs. 77.15%), showcasing the system’s proficiency at navigating intricate reasoning tasks with enhanced cost-efficiency.